OpenStudy (anonymous):

How do you simplify (x/x^2-y^2)+(y/x^2-y^2) ?

OpenStudy (anonymous):

i think you add like terms.

OpenStudy (anonymous):

\[x/x ^{2} - y ^{2} + y/x ^{2} - y ^{2}\]\[x/x ^{2} + y/x ^{2}- y ^{2} - y ^{2}\]\[(x+y)/x ^{2} -2y ^{2}\]

OpenStudy (anonymous):

@rachaellla is correct, in this case you have the same denominator (x^2-y^), so you simply add the numerator and the result should be like @stargirlzilin has done above

OpenStudy (anonymous):

@stargirlzilin That's what I did/got, but it isn't one of the choices

OpenStudy (anonymous):

@stargirlzilin is that 2 in your answers denominator supposed to be there?

OpenStudy (anonymous):

WAIT OH I typed the question wrong it's - not +

OpenStudy (anonymous):

@snowflakepixie remember the denominator is a difference of two squares: ie \[x ^{2}-y ^{2} = (x+y)(x-y)\], now can you simplify further?

OpenStudy (anonymous):

GAH I'M CONFUSING MYSELF SO MUCH

OpenStudy (anonymous):

\[(x+y) / (x ^{2}- y ^{2}) = (x+y) / (x+y)(x-y) = 1/(x-y)\]

OpenStudy (anonymous):

Alright alright one minute... it's